OpenStudy (anonymous):

a.Investing is an important topic for individuals in all areas. Suppose you wish to invest $10,000 in a money market account that earns 4% interest compounded annually. How much would the investment be worth after 10 years? Show all calculations. b.Now, suppose that the above investment earns interest compounded quarterly. How much would the same investment be worth after 10 years? Why does the result differ from the previous one?

OpenStudy (anonymous):

\[10,000\times (1.04)^{10}\] for the first one

OpenStudy (anonymous):

wow thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wow the second one is wrong! it is \[10,000\times (1+\frac{.04}{12})^{120}\]